You are thinking of investing in Nikki T's, Inc. You have only the following information on the firm at year-end 2018: net income is $180,000, total debt is $2.80 million, and debt ratio is 60 percent. What is Nikki T's ROE for 2018? (Do not round intermediate calculations and round your final answer to 2 decimal places.)
Answer #1

debt ratio=debt/total assets

total assets=(2.8/0.6) million

=$4.67 million

total assets=debt+equity

equity=(4.67-2.8)=$1.87 million(Approx)

=$1866666.67

ROE=net income/equity

=180,000/1866666.67

=9.64%(Approx).